i dont understand..
-54x-60y=6
-72x-80y=8

Answer by w_parminder(53)

-54x-60y=6
taking 6 common, we get
6(-9x-10y)=6
OR -9x-10y=1 ....... eq (1)

-72x-80y=8
taking 8 common, we get
8(-9x-10y)=8
OR -9x-10y=1 ....... eq (2)
Now as equation (1) and (2) are exactly same, we can just say that these pair of equations have infinitely many solutions. These equations can be solved separately but not this pair of equations.